Find dy/dx and d²y/dx² without eliminating the parameter.


a.) x= e^(2t) , y= 1+cos(t)

b.) x= acosh(t) , y= bsinh(t)


1
Expert's answer
2021-01-18T06:08:40-0500

Solution

a.)   x= e^(2t) , y= 1+cos(t) =>

dy/dx=(dy/dt)/(dx/dt)= (-sin(t))/(2e2t)=-e-2t*sin(t)/2

d2y/dx2=(d(dy/dx)/dt)/(dx/dt)=(e-2t *sin(t)- e-2t *cos(t)/2)/(2e2t)= e-4t (2*sin(t)-cos(t))/4 

b.)   x= acosh(t) , y= bsinh(t) =>

dy/dx=(dy/dt)/(dx/dt)=(b/a)*cosh(t)/sinh(t)=(b/a)*coth(t)

d2y/dx2=(d(dy/dx)/dt)/(dx/dt)=(b/a)*(-cosech(t)2)/(a*sinh(t)) = -b/(a2*sinh(t)3)

 Answer

a.)  dy/dx=-e-2t*sin(t)/2, d2y/dx2=e-4t (2*sin(t)-cos(t))/4 

b.)  dy/dx=(b/a)*coth(t), d2y/dx2= -b/(a2*sinh(t)3)